The annual precipitation amounts in a certain mountain range are normally distributed with a mean of 107 inches, and a standard deviation of 12 inches. What is the probability that the mean annual precipitation during 36 randomly picked years will be less than 109.8 inches?
Answer by stanbon(75887)

The annual precipitation amounts in a certain mountain range are normally distributed with a mean of 107 inches, and a standard deviation of 12 inches. What is the probability that the mean annual precipitation during 36 randomly picked years will be less than 109.8 inches?
------------------------------
z(109.8) = (109.8-107)/sqrt(36) = 0.4667
P(x-bar < 109.8) = P(z < 0.4667) = 0.6796
===========
Cheers,
Stan H.
===========